Accounting for Bonds Sold at a Discount
The Biltmore National Bank raised capital through the sale of $3 million face value of 8% coupon rate, 10-year bonds.

The bonds paid interest semiannually and were sold at a time when equivalent risk-rated bonds carried a yield rate of 10%.

Calculate the proceeds that The Biltmore National Bank received from the sale of the 8% bonds.
Round your answer to the nearest dollar.

$Answer


How will the bonds be disclosed on Biltmore's balance sheet immediately following the sale?
Round your answers to the nearest dollar.

Balance sheet disclosure (following sale):

Bonds payable Answer
Less: Bonds discount (enter as negative) Answer
Bonds payable(net) Answer


Calculate the interest expense on the bonds for the first year that the bonds are outstanding.
Do not round until final answer. Round answers to the nearest dollar.

First six months Answer
Second six months Answer


Calculate the book value of the bonds at the end of the first year.
Do not round until final answer. Round answer to the nearest dollar.

$Answer
Answer #1

Solution

Biltmore National Bank

  1. Calculation of the proceeds that the Biltmore National Bank received from sale of the 8% bonds:

Bond issue proceeds = present value of bond + present value of interest payments

Present value of bonds = 3,000,000 x (P/F, 5%, 20)

Period, n = 10 x 2 = 20

Semiannual yield rate = 10%/2 = 5%

PV of bonds = 3,000,000 x 0.3769 = $1,130,700

PV of interest payments –

Semiannual interest = 3,000,000 x 8% x 6/12 = $120,000

PV of interest = 120,000 x (P/A, 5%, 20)

= 120,000 x 12.462 = $1,495,440

Bond issue price = 1,130,700 + 1,495, 440 = $2,626,140

Discount on bonds payable = 3,000,000 – 2,626,140 = $373,860

  1. Balance sheet disclosure immediately after sale:

Balance sheet disclosure (following sale):

Bonds Payable

$3,000,000

Less: Bonds discount

($373,860)

Bonds Payable (net)

$2,626,140

  1. Calculation of the interest expense on the bonds for the first year that the bonds are outstanding:

First six months –

Interest expense = carrying value of bonds x market yield rate

= 2,626,140 x 5% = $131,307

Second six months = (2,626,140 + discount amortization) x 5%

Discount amortization = 131,307 – 120,000 = 11,307

Carrying value of bond = 2,626,140 + 11,307 = $2,627,447

Interest expense = 2,627,447 x 5% = $131,872

Interest Expense

First six months

$131,307

Second six months

$131,872

  1. Book value of bonds at end of first year:

Book value after first interest payment = 2,627,447 (computed above)

Book value after second interest payment –

Discount amortization = 131,872 – 120,000 = 11,872

Book value = 2,627,447 + 11,872 = $2,639,319

Book value of bonds at the end of first year = 2,639,319